### v2

@@ Abstract
With a view on bilevel and PDE-constrained optimisation, we develop iterative estimates $\widetilde{F^{\prime}}(x^{k})$ of $F^{\prime}(x^{k})$ for composite functions $F:=J\circ S$, where $S$ is the solution mapping of the inner optimisation problem or PDE. The idea is to form a single-loop method by interweaving updates of the iterate $x^{k}$ by an outer optimisation method, with updates of the estimate by single steps of standard optimisation methods and linear system solvers. When the inner methods satisfy simple tracking inequalities, the differential estimates can almost directly be employed in standard convergence proofs for general forward-backward type methods. We adapt those proofs to a general inexact setting in normed spaces, that, besides our differential estimates, also covers mismatched adjoints and unreachable optimality conditions in measure spaces. As a side product of these efforts, we provide improved convergence results for nonconvex Primal-Dual Proximal Splitting (PDPS).
@@ 1 Introduction
First-order methods are slow. To be precise, they require a high number of iterations, but if those iterations are fast, they have the chance to practically overpower second-order methods with expensive iterations. In PDE-constrained or bilevel optimisation, the steps of basic first-order methods are very expensive, involving the solution of the inner problem or PDE and its adjoint. To make first-order methods fast, it is, therefore, imperative to reduce the cost of solving these subproblems—for instance, by employing inexact solution schemes.
Consequently, especially in the machine learning community, an interest has surfaced in single-loop methods for bilevel optimisation; see [28] and references therein. Many of these methods are very specific constructions. In [16] we started work on a more general approach to PDE-constrained optimisation: we showed that on each step of an outer primal-dual optimisation method, we can take single steps of standard linear system splitting schemes for the PDE constraint and its adjoint, and still obtain a convergent method that is computationally significantly faster than solving the PDEs exactly. In [29] we then presented an approach to bilevel optimisation that allowed general inner and adjoint algorithms that satisfy certain tracking inequalities. These were proved for standard splitting schemes for the adjoint equation, and for forward-backward splitting and the Primal-Dual Proximal Splitting (PDPS) of [4] for the inner problem. The overall analysis was still tied to bilevel optimisation in Hilbert spaces, with forward-backward splitting as the outer optimisation method.
Writing $F=J\circ S_{u}$ for a solution mapping $S_{u}:X\to U$ of an inner problem (PDE or an optimisation problem, or multiple nested levels of such), and a differentiable function $J:U\to\mathbb{R}$, on normed spaces $X$ and $U$, we are interested in the solution of composite optimisation problems of the form
(1) $\min_{x\in X}\leavevmode\nobreak\ F(x)+G(x),$
@@
for $G$ convex but possibly nonsmooth, and $\Xi\in\mathbb{L}(X;X^{*})$ skew-adjoint. If $\Xi=0$, then this optimality condition is typically necessary for (1). More generally, the operator allows the modelling of primal-dual problems, and treating the PDPS and Douglas–Rachford splitting as generalised forward-backward splitting methods [7].
Our contributions are as follows. In Section 2, which forms our inner theory,
@@
    we show in general normed spaces that we can approximate in a single-loop fashion the differentials of compositions $F=J\circ S_{u}$, given abstract inner and adjoint algorithms for $S_{u}$, satisfying certain tracking inequalities.
In contrast to [29] and, indeed, all single-loop bilevel optimisation methods that we are aware of, our approach can also work with the adjoint dimension reduction trick typically employed in PDE-constrained optimisation. We show that, subject to additive error terms with a bounded sum, the differential estimates $\widetilde{F^{\prime}}(x^{k})$ satisfy standard smoothness properties, such as Lipschitz differential and the two- and three-point descent inequalities [31, 7]. Based on this, in Section 4, which forms our outer theory,
@@
    we then prove various forms of convergence of general inexact splitting methods for (2).
In Section 4.1 we also elaborate on the aforementioned connection of primal-dual methods to Eq. 2. To facilitate the analysis of such methods, we first introduce in Section 3 operator-relative variants of the descent inequality.
Through our approach to inexactness, besides gradient estimates for multilevel problems, we can model mismatched adjoints [17], and difficult-to-solve-exactly optimality conditions in measure spaces [34]. We also adopt the approach of [34] to optimisation in normed spaces: instead of Bregman divergences, we construct an inner product structure with a self-adjoint $M\in\mathbb{L}(X;X^{*})$. Our work is related to the study of gradient oracles for smooth convex optimisation in [8], and for nonconvex composite optimisation in [10, 20], both in finite-dimensional Euclidean spaces. Based on sufficient descent and the Kurdyka–Łojasiewicz property, [22] also study inexact methods in $\mathbb{R}^{n}$. Moreover, [3] introduce approaches to control model inexactness in proximal trust region methods, and [27] in non-single-loop gradient methods for bilevel optimisation.
Not content to merely adapt existing proofs to inexact steps and normed spaces, we also present some improvements, especially for the nonconvex PDPS of [30]. We do, however, treat a slightly simplified problem. The original PDPS of [4] applies to $\min_{x}g(x)+h(K(x))$ with $K$ linear and $g$ and $h$ convex. The extension of [30] allows $K$ to be nonlinear. It is further analysed in [5, 6, 18, 11], with a simplified review of the different variants in Banach space in [32]. An alternative extension in [19] allows $g$ and $h$ to be semiconvex. We, instead, consider the special case $\min_{x}f(x)+g(x)+h(Kx)$ with $K$ linear, but $f$ nonconvex, and
@@
We write $\mathbb{L}(X;Y)$ for the space of bounded linear operators between the normed spaces $X$ and $Y$, and $\operatorname{Id}$ for the identity operator. $X^{*}$ stands for the dual space of $X$. When $X$ is Hilbert, we identify $X^{*}$ with $X$. We write $\langle x,y\rangle$ for an inner product, $\langle x^{*}|x\rangle_{X^{*},X}$ for a dual product. We call $M\in\mathbb{L}(X;X^{*})$ self-adjoint if the restriction $M^{*}|X=M$, and positive semi-definite if $\langle x|Mx\rangle_{X,X^{*}}\geq 0$ for all $x\in X$. If both hold, we set $\|x\|_{M}:=\sqrt{\langle Mx|x\rangle}$, and write $\mathbb{O}_{M}(x,r)$ for the radius-$r$ open ball at $x$ in the $M$-(semi-)norm. We also write $M\geq N$ if $M-N$ is positive semi-definite. We extensively use the vectorial Young’s inequality
@@
2 Tracking estimate recursion
Let $J:U\to\mathbb{R}$ and $S_{u}:X\to U$ be Fréchet differentiable on normed spaces $X$ and $U$. We consider the functional
@@ 2 Tracking estimate recursion
As $S_{u}$ and its differential can be expensive to compute, given an iterate $x^{k}$ of an arbitrary outer algorithm for minimising an objective that involves $F$, such as (1), we estimate $S_{u}(x^{k})$ by $u^{k+1}\in U$, and $S_{u}^{\prime}(x^{k})$ by $p^{k+1}\in\mathbb{L}(X;U)$, that is, we estimate
@@
When $X$ is Hilbert, we write $\widetilde{\nabla F}(x^{k})$ for the Riesz representation of $\widetilde{F^{\prime}}(x^{k})$. We do not provide a single explicit formula for $u^{k+1}$ and $p^{k+1}$, instead, we assume them to satisfy tracking estimates as in [16, 29]. We formulate these tracking estimates—that are essentially contractivity estimates with suitable penalties for parameter change—in Section 2.1, and provide examples of inner and adjoint methods that satisfy them. We follow with an estimate on the error of approximation of $F^{\prime}(x^{k})$ by $\widetilde{F^{\prime}}(x^{k})$ in Section 2.2. Our goal is to derive, in Section 2.3, variants of standard descent inequalities and Lipschitz bounds for the estimate $\widetilde{F^{\prime}}(x^{k})$.
@@
Suppose $S_{u}(x)$ arises from the satisfaction of
(3) $0=T(S_{u}(x),x)\quad\text{for a}\quad T:U\times X\to W_{*}\quad\text{with}\quad W_{*}\text{ a normed space},$
modelling, e.g., a PDE or the first-order optimality conditions of an inner optimisation problem, both parametrised by $x$. By implicit differentiation, subject to sufficient differentiability and (3) holding in a neighbourhood of $x$, we obtain the basic adjoint
@@
where $S_{u}^{\prime}(x)\in\mathbb{L}(X;U)$, $T^{(u)}(S_{u}(x),x)\in\mathbb{L}(U;W)$, and $T^{(x)}(S_{u}(x),x)\in\mathbb{L}(X;W)$. Hence, following the derivation of adjoint PDEs in, e.g., [14, §1.6.2] or [24, §1.2], assuming $T^{(u)}(S_{u}(x),x)$ to be invertible, we solve from (4) that
$[J\circ S_{u}]^{\prime}(x)=J^{\prime}(S_{u}(x))S_{u}^{\prime}(x)=w_{x}T^{(x)}(S_{u}(x),x),$
for a $w_{x}=S_{w}(x)\in W$ satisfying the reduced adjoint
(5) $w_{x}T^{(u)}(S_{u}(x),x)+J^{\prime}(S_{u}(x))=0.$
@@
2.1 Basic constructions and assumptions
To track the inexact computations of inner and adjoint variables across iterations, we introduce abstract but verifiable conditions that quantify how closely the computed values follow the outputs of the exact inner and adjoint solution mappings evaluated at the current outer iterate. These tracking assumptions ensure that the accumulated errors remain controlled and that the approximate gradient remains meaningful for descent. The following assumption formalises this idea. There, and in the rest of this section, unless otherwise explicitly mentioned, the distances $b_{X}$, $d_{X^{*}}$, $d_{U}$, and $d_{W}$ are entirely formal expressions. Typically they will be given by norms, but the theory does not require this. Likewise, the “target” $F^{\prime}(x^{k})$ will typically be the Fréchet derivative of $J\circ S_{u}$ at $x^{k}$, but this need not be the case; it is likewise an entirely formal expression. We write squared distances as $b_{X}^{2}(x,y):=b_{X}(x,y)^{2}$.

### v3

@@ Abstract
With a view on bilevel and PDE-constrained optimisation, we develop iterative estimates $\widetilde{F^{\prime}}(x^{k})$ of $F^{\prime}(x^{k})$ for composite functions $F:=J\circ S$, where $S$ is the solution mapping of the inner optimisation problem or PDE, the latter seen in this work as a special case of the former.. The idea is to form a single-loop method by interweaving updates of the iterate $x^{k}$ by an outer optimisation method, with updates of the estimate by single steps of standard optimisation methods and linear system solvers. When the inner methods satisfy simple tracking inequalities, the differential estimates can almost directly be employed in standard convergence proofs for general forward-backward type methods. We adapt those proofs to a general inexact setting in normed spaces, that, besides our differential estimates, also covers mismatched adjoints and unreachable optimality conditions in measure spaces. As a side product of these efforts, we provide improved convergence results for nonconvex Primal-Dual Proximal Splitting (PDPS).
@@ 1 Introduction
First-order methods are slow. To be precise, they require a high number of iterations, but if those iterations are fast, they have the chance to practically overpower second-order methods with expensive iterations. In bilevel or PDE-constrained optimisation—the latter seen in this work as a special case of the former—the steps of basic first-order methods are very expensive, involving the solution of the inner problem or PDE and its adjoint. To make first-order methods fast, it is, therefore, imperative to reduce the cost of solving these subproblems—for instance, by employing inexact solution schemes.
Consequently, especially in the machine learning community, an interest has surfaced in single-loop methods for bilevel optimisation; see [29] and references therein. Many of these methods are very specific constructions. In [18] we started work on a more general approach to PDE-constrained optimisation: we showed that on each step of an outer primal-dual optimisation method, we can take single steps of standard linear system splitting schemes for the PDE constraint and its adjoint, and still obtain a convergent method that is computationally significantly faster than solving the PDEs exactly. The adjoint equation is used to compute differentials of the PDE or inner problem solutions with respect to its parameters. In [30] we then presented an approach to bilevel optimisation that allowed general inner and adjoint algorithms that satisfy certain tracking inequalities. These were proved for standard splitting schemes for the adjoint equation, and for forward-backward splitting and the Primal-Dual Proximal Splitting (PDPS) of [4] for the inner problem. The overall analysis was still tied to bilevel optimisation in Hilbert spaces, with forward-backward splitting as the outer optimisation method.
Our goal, in this work, is to develop a general theory of optimisation methods for bilevel and, by extension, multilevel problems. This theory is based on the idea of algorithmic single-loop differential estimation, with a flexible choice of outer, inner, and adjoint algorithms. The inner and adjoint methods only need to satisfy the aforementioned tracking inequalities. The outer method only has to tolerate controlled inexactness of differentials involving the inner solution mapping.
Problem setup
Write
$F=J\circ S_{u}$
for a mapping $S_{u}:X\to U$ and a differentiable function $J:U\to\mathbb{R}$, on normed spaces $X$ and $U$. Typically, but not necessarily, $S_{u}(x)$ is an inner solution mapping that arises from the satisfaction of
(1) $0=T(S_{u}(x),x)\quad\text{for a}\quad T:U\times X\to W_{*}\quad\text{with}\quad W_{*}\text{ a normed space}.$
The idea is that $T$ encodes the optimality conditions of an inner problem, parametrised by $x$, or a PDE, likewise parametrised by $x$. We are then interested in the solution of composite optimisation problems of the form
(2) $\min_{x\in X}\penalty 10000\ F(x)+G(x),$
@@
for $G$ convex but possibly nonsmooth, and $\Xi\in\mathbb{L}(X;X^{*})$ skew-adjoint. If $\Xi=0$, then this optimality condition is typically necessary for (2). More generally, the operator allows the modelling of primal-dual problems, and treating the PDPS and Douglas–Rachford splitting as generalised forward-backward splitting methods [7, 32]. We will discuss such formulations in more detail in Section˜4.
Contributions
Our contributions are as follows. In Sections˜2 and 3, which form our inner theory,
@@
    we show in general normed spaces that we can approximate in a single-loop fashion the differentials of compositions $F=J\circ S_{u}$, given abstract inner and adjoint algorithms for $S_{u}$, satisfying certain tracking inequalities. We then illustrate how standard algorithms satisfy these inequalities.
The corresponding results on sequences of scalars, on which these results are based on, are relegated to Appendix˜A. In contrast to [30] and, indeed, all single-loop bilevel optimisation methods that we are aware of, our approach can also work with the adjoint dimension reduction trick typically employed in PDE-constrained optimisation. We show that, subject to additive error terms with a bounded sum, the differential estimates $\widetilde{F^{\prime}}(x^{k})$ satisfy standard smoothness properties, such as Lipschitz differential and the two- and three-point descent inequalities [32, 7]. Based on this, in Section˜4, which forms our outer theory,
@@
    we prove various forms of convergence of general inexact splitting methods for (3).
In Section˜4.1 we also elaborate on the aforementioned connection of primal-dual methods to Eq.˜3. To facilitate the analysis of such methods, we introduce in Section˜5 operator-relative variants of the descent inequality. We then interpret the conditions of Section˜4 for outer forward-backward splitting and outer PDPS in Section˜6.
Not content to merely adapt existing proofs to inexact steps and normed spaces, we also present some improvements, especially for the nonconvex PDPS of [31]. We do, however, treat a slightly simplified problem. The original PDPS of [4] applies to $\min_{x}g(x)+h(K(x))$ with $K$ linear and $g$ and $h$ convex. The extension of [31] allows $K$ to be nonlinear. It is further analysed in [5, 6, 20, 12], with a simplified review of the different variants in Banach space in [33]. An alternative extension in [21] allows $g$ and $h$ to be semiconvex. We, instead, consider the special case $\min_{x}f(x)+g(x)+h(Kx)$ with $K$ linear, but $f$ nonconvex, and, in Section˜6,
@@
Examples
The following examples illustrate the kind of algorithms that can be constructed, and whose convergence can be proved using our general theory. Our overall theory is, however, much more general than these two specific algorithms, as it allows the overall method to be composed of arbitrary inner, adjoint, and outer algorithms. Instead of gradients in Hilbert spaces, we will, moreover, work with differentials in general normed spaces.
Example 1.1 (Forward-backward–linear system splitting–forward-backward).
Consider the bilevel optimisation problem
$\min_{x,u}J(u)+G(x)\quad\text{subject to}\quad u\in\operatorname*{arg\,min}_{x}f(u;x)+g(u;x).$
Assume for simplicity that $f$ and $g$ are twice differentiable, convex in the first variable, and all the spaces are Hilbert. Then we can rewrite the problem as Eqs.˜2 and 1 by setting
$T(u,x)=\nabla_{u}f(u;x)+\nabla_{u}g(u;x).$
A standard forward backward method with step length parameter $\tau>0$ would update
$x^{k+1}:=\operatorname{prox}_{\tau G}(x^{k}-\nabla F(x^{k})),$
however the computation of $\nabla F(x)=\nabla S_{u}(x)^{*}\nabla J(S_{u}(x))$ is generally expensive. We therefore propose to form the iteratively updated single-loop differential estimate $\widetilde{\nabla F}(x^{k})$, as in the overall algorithm
(4) $\left\{\begin{aligned} u^{k+1}&:=\operatorname{prox}_{\sigma\nabla g(\,\boldsymbol{\cdot}\,;x^{k})}(u^{k}-\sigma\nabla f(u^{k};x^{k})),&\text{(inner step)}\\ w^{k+1}&:=-N^{-1}(Mw^{k}+\nabla J(u^{k+1}))\quad\text{where}\quad N+M=\nabla_{u}T(u^{k+1},x^{k}),&\text{(adjoint step)}\\ \widetilde{\nabla F}(x^{k})&:=\nabla_{x}T(u^{k+1},x^{k})w^{k+1},&\text{(diff.\penalty 10000\ transform)}\\ x^{k+1}&:=\operatorname{prox}_{\tau G}(x^{k}-\widetilde{\nabla F}(x^{k})).&\text{(outer step)}\end{aligned}\right$.
The first line is a standard forward-backward step for the inner variable. The second line applies, e.g., Gauss–Seidel or block-Gauss–Seidel splitting (depending on the choice of the operators $N$ and $M$) to the linear equation
$\nabla_{u}T(u^{k+1},x^{k})w^{k+1}+\nabla J(u^{k+1})=0.$
Of course, the choice $M=0$ and $N=\nabla_{u}T(u^{k+1},x^{k})$ would solve this equation exactly. The third line of (4) transforms the inner and adjoint iterates $u^{k+1}$ and $w^{k+1}$ into the differential estimate $\widetilde{\nabla F}(x^{k})$. As we later elaborate in Example˜2.1, this construction is motivated by the reduced adjoint equation
$\nabla_{u}T(S_{u}(x),x)w_{x}+\nabla J(S_{u}(x))=0\quad\text{whose solution satisfies}\quad\nabla F(x)=\nabla_{x}T(S_{u}(x),x)w_{x}.$
We treat specific inner and adjoint methods and differential transformations in Section˜3 after the general theory in Section˜2. The fourth line of (4) is an outer inexact forward-backward step. We present a general theory of outer/inexact methods in Section˜4, and provide examples in Section˜6.
Example 1.2 (A simple PDE-constrained optimisation problem).
With $U\subset H^{1}(\Omega)$ and $X\subset L^{2}(\Omega)$, define for admissible $x\in X$ the linear operator $A_{x}\in\mathbb{L}(U;U^{*})$ by $A_{x}u:=\nabla^{*}[x\cdot\nabla u]$, and consider the PDE-constrained optimisation problem
$\min_{x,u}J(u)+G(x)\quad\text{subject to}\quad A_{x}u=b,$
where we use $G$ to restrict $x$ to a domain where the PDE is well-posed. Setting $T(u,x)=A_{x}u-b$, we are again the setup of Eqs.˜2 and 1. Observing that $\nabla_{u}T(u^{k+1},x^{k})=A_{x}$ and $\nabla_{x}T(u^{k+1},x^{k})w^{k+1}=\nabla u^{k+1}\cdot\nabla w^{k+1}$, similarly to Example˜1.1, but this time also applying a linear system splitting scheme to the PDE itself, we arrive at the algorithm
$\left\{\begin{aligned} \text{split }&N+M:=A_{x^{k}},&\text{(operator splitting)}\\ u^{k+1}&:=-N^{-1}(Mu^{k}-b),&\text{(PDE step)}\\ w^{k+1}&:=-N^{-1}(Mw^{k}+\nabla J(u^{k+1})),&\text{(adjoint step)}\\ \widetilde{\nabla F}(x^{k})&:=\nabla u^{k+1}\cdot\nabla w^{k+1},&\text{(diff.\penalty 10000\ transform)}\\ x^{k+1}&:=\operatorname{prox}_{\tau G}(x^{k}-\widetilde{\nabla F}(x^{k})).&\text{(outer step)}\end{aligned}\right$.
Note that, to use standard schemes such as Gauss–Seidel, we need to work in finite-dimensional subspaces. However, in infinite dimensions, the abstract splitting $A_{x}=N+M$ can applied on the block structure of $A_{x}$ in more complex problems.
Further related work
Through our general approach to inexactness in Section˜4, independent of the differential estimation theory of Section˜2, besides differential estimates for multilevel problems, we can model mismatched adjoints [19], and difficult-to-solve-exactly optimality conditions in measure spaces [35]. We also adopt the approach of [35] to optimisation in normed spaces: instead of Bregman divergences, we construct an inner product structure with a self-adjoint $M\in\mathbb{L}(X;X^{*})$. Our work is related to the study of gradient oracles for smooth convex optimisation in [9], and for nonconvex composite optimisation in [11, 22], both in finite-dimensional Euclidean spaces. Based on sufficient descent and the Kurdyka–Łojasiewicz property, [24] also study inexact methods in $\mathbb{R}^{n}$. Moreover, [3] introduce approaches to control model inexactness in proximal trust region methods, and [28] in non-single-loop gradient methods for bilevel optimisation.
@@
We write $\mathbb{L}(X;Y)$ for the space of bounded linear operators between the normed spaces $X$ and $Y$, and $\operatorname{Id}$ for the identity operator. $X^{*}$ stands for the dual space of $X$. When $X$ is Hilbert, we identify $X^{*}$ with $X$. We write $\langle x,y\rangle$ for an inner product, $\langle x^{*}|x\rangle_{X^{*},X}$ for a dual product. We call $M\in\mathbb{L}(X;X^{*})$ self-adjoint if the restriction $M^{*}|X=M$, and positive semi-definite if $\langle x|Mx\rangle_{X,X^{*}}\geq 0$ for all $x\in X$. If both hold, we set $\|x\|_{M}:=\sqrt{\langle Mx|x\rangle}$, and write $\mathbb{O}_{M}(x,r)$ for the radius-$r$ open ball at $x$ in the MM-(semi-)norm. The open ball in the standard norm of the space is denoted by $\mathbb{O}(x,r)$. We also write $M\geq N$ if MNM-N is positive semi-definite. We extensively use the vectorial Young’s inequality
@@
For sets $A$ and $B$, we often write $\langle A-B|x\rangle\geq 0$, which means that $\langle y-z|x\rangle\geq 0$ for all $y\in A$ and $z\in B$.
@@
2 Differential estimation from tracking inequalities
Let $J:U\to\mathbb{R}$ and $S_{u}:X\to U$ be Fréchet differentiable on normed spaces $X$ and $U$. We consider
@@ 2 Differential estimation from tracking inequalities
Typically, but not necessarily in the overall theory, $S_{u}(x)$ arises from the satisfaction of (1).
As $S_{u}$ and its differential can be expensive to compute, given an iterate $x^{k}$ of an arbitrary outer algorithm for minimising an objective that involves $F$, such as (2), we estimate $S_{u}(x^{k})$ by an inner iterate $u^{k+1}\in U$, and $S_{u}^{\prime}(x^{k})$ by an adjoint iterate $p^{k+1}\in\mathbb{L}(X;U)$, that is, we estimate
@@
When $X$ is a Hilbert space, we write $\widetilde{\nabla F}(x^{k})$ for the Riesz representation of $\widetilde{F^{\prime}}(x^{k})$. We do not provide a single explicit formula for $u^{k+1}$ and $p^{k+1}$. Instead, we assume them to satisfy tracking estimates as in [18, 30]. We formulate these tracking estimates—that are essentially contractivity estimates with suitable penalties for parameter change—in Section˜2.1. Our goal is to derive, in Section˜2.2, variants of standard descent inequalities and Lipschitz bounds for the estimate $\widetilde{F^{\prime}}(x^{k})$. In Section˜3 we provide examples of inner and adjoint methods that satisfy the tracking inequalities.
@@
Suppose $S_{u}(x)$ arises from the satisfaction of (1). By implicit differentiation, subject to sufficient differentiability and (1) holding in a neighbourhood of $x$, we obtain the basic adjoint
@@
where $S_{u}^{\prime}(x)\in\mathbb{L}(X;U)$, $T^{(u)}(S_{u}(x),x)\in\mathbb{L}(U;W)$, and $T^{(x)}(S_{u}(x),x)\in\mathbb{L}(X;W)$. Hence, following the derivation of adjoint PDEs in, e.g., [16, §1.6.2] or [8, §1.2], assuming $T^{(u)}(S_{u}(x),x)$ to be invertible, we solve from (5) that
$[J\circ S_{u}]^{\prime}(x)=J^{\prime}(S_{u}(x))S_{u}^{\prime}(x)=S_{w}(x)T^{(x)}(S_{u}(x),x),$
for a $S_{w}(x)\in W$ satisfying the reduced adjoint
(6) $S_{w}(x)T^{(u)}(S_{u}(x),x)+J^{\prime}(S_{u}(x))=0.$
@@
2.1 The tracking assumption
To track the inexact computations of inner and adjoint variables across iterations, we introduce verifiable conditions that quantify how closely the computed values follow the outputs of the exact inner and adjoint solution mappings evaluated at the current outer iterate. These tracking assumptions ensure that the accumulated errors remain controlled and that the approximate gradient remains meaningful for descent. The following assumption formalises this idea.